A 23 kg suitcase is being pulled by its handle at a constant speed across level ground. The suitcase handle is at an angle of 25 degrees above the horizontal. If the normal force exerted on the suitcase is 180 N, what is the magnitude of the force being applied to the handle?
Answer #1

Here N is the normal force on the suitcase , F is the applied force and \theta = 25
taking equilibrium in the vertical direction
FSin25 + N = mg
FSin25 + 180 = 23*9.81
F = 107.969 N
Hence this will be the applied force